OpenStudy (anonymous):

A video game sets the points needed to reach the next level based on the function g(x) = 8(2)^x + 1, where x is the current level. The hardest setting promises to multiply the points needed in each level according to the function h(x) = 3x. How many points will a player need on the hardest setting of level 5?

OpenStudy (kirbykirby):

is it \( \large 8(2)^{x+1} \) or \(\large 8(2)^x + 1\)

OpenStudy (anonymous):

the first one

OpenStudy (kirbykirby):

so the number of points at level 5 would be: \(\large g(5) = 8(2)^{5+1}=8(2)^6=512\) But, it says that it also multiplies the points at each level by the function \(h(x)=3x\) So, \(h(5)=3(5)=15\) Thus, it will multiply 512 by 15 = 7680

OpenStudy (anonymous):

thank you

OpenStudy (kirbykirby):

:)